# Towards Adversarially Robust Dataset Distillation by Curvature Regularization

### v1

@@ 4.1 Formulation of the Robust Dataset Distillation Problem
4.2 Theoretical Bound of Robustness
Our aim is to create a method that allows us to efficiently and reliably introduce robustness into distilled datasets, thus we will start by exploring the theoretical connections between dataset distillation and adversarial robustness.
Conveniently, previous works (Jetley et al., 2018; Fawzi et al., 2018) have studied the adversarial robustness of neural networks via the geometry of the loss landscape. Here we find connections between standard training procedures and dataset distillation to provide a theoretical bound for the adversarial loss of models trained with distilled datasets.
@@ 4.2 Theoretical Bound of Robustness
$\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\tilde{\ell}_{\rho}^{adv}(\mathbf{x})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|{\nabla\ell(\mathbf{x})}\|+\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x}),$ (6)
where $\lambda_{1}$ is the largest eigenvalue of the Hessian matrix $\mathbf{H(\ell(x))}$.
Then, we have the proposition:
@@
~$\tilde{\ell}_{\rho}^{adv}(\mathbf{x}^{\prime})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|\nabla\ell(\mathbf{x})\|+\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x})+L\sigma.$ (7)
Given the assumption of convexity in the loss function, we can further observe that in a convex landscape the gradient magnitude tends to be lower, particularly near the optimal points. Therefore, in the context of a convex loss function and a well-distilled dataset, the gradient term $\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|\nabla\ell(\mathbf{x})\|$ contribute insignificantly to the overall value of the inequality. This insignificance is amplified by the presence of the curvature term, $\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x})$, which provides a sufficient descriptor of the loss landscape under our assumptions.
Hence, it is reasonable to simplify the expression by omitting the gradient term, resulting in a focus on the curvature term, which is more representative of the convexity assumption and the characteristics of a well-distilled dataset. The revised expression would then be:
@@
In Appendix A, we give a more thorough proof of the proposition and discuss the validity of some of the assumptions made.
4.3 Geometric Regularization for Adversarial Robust Dataset
Based on our theoretical discussion in Section 4.2, we propose a method, GUARD (Geometric regUlarization for Adversarial Robust Dataset). Since the theorem suggests that the upper bound of the adversarial loss is mainly determined by the curvature of the loss function, we modify the distillation process so that the trained model has a loss function with a low curvature with respect to real data.
To reduce $\lambda_{1}$ in Eq. 8 requires computing the Hessian matrix and get the largest eigenvalue $\lambda_{1}$, which is quite computationally expensive. Here we find an efficient approximation of it. Let $\mathbf{v_{1}}$ be the unit eigenvector corresponding to $\lambda_{1}$, then the Hessian-vector product is
@@ 4.3 Geometric Regularization for Adversarial Robust Dataset
Previous works (Fawzi et al., 2018; Jetley et al., 2018; Moosavi-Dezfooli et al., 2019) have empirically shown that the direction of the gradient has a large cosine similarity with the direction of $\mathbf{v_{1}}$ in the input space of neural networks. Instead of calculating $\mathbf{v_{1}}$ directly, it is more efficient to take the gradient direction as a surrogate of $\mathbf{v_{1}}$ to perturb the input $\mathbf{x}$. So we replace the $\mathbf{v}_{1}$ above with the normalized gradient $\mathbf{z}=\frac{\nabla\ell(\mathbf{x}))}{\|\nabla\ell(\mathbf{x}))\|}$, and define the regularized loss $\ell_{R}$ to encourage linearity in the input space:
@@ 4.4 Engineering Specification
5 Experiments
5.1 Experiment Settings
@@
For a systematic evaluation of our method, we investigate the top-$1$ classification accuracy of models trained on data distilled from three datasets: ImageNette (Howard, 2018), Tiny ImageNet (Le and Yang, 2015), and ImageNet (Deng et al., 2009). ImageNette is a subset of ImageNet containing 10 easy-to-classify classes. Tiny ImageNet is a scaled-down subset of ImageNet, containing 200 classes and 100,000 downsized 64x64 images. We train networks using the distilled datasets and subsequently verify the network’s performance on the original datasets. For consistency in our experiments across all datasets, we use the standard ResNet18 architecture (He et al., 2016) to synthesize the distilled datasets and evaluate their performance. During the squeeze step of the distillation process, we train the model on the original dataset over 50 epochs using a learning rate of 0.025. Based on preliminary experiments, we determined that the settings $h=3$ and $\lambda=100$ provide an optimal configuration for our regularizer. In the recover step, we perform 2000 iterations to synthesize the images and run 300 epochs to generate the soft labels to obtain the full distilled dataset. In the evaluation phase, we train a ResNet18 model on the distilled dataset for 300 epochs, before assessing it on the test set of the original dataset.
5.2 Comparison with Other Methods
@@
As of now, there is only a small number of dataset distillation methods that can achieve good performance on ImageNet-level datasets, therefore our choices for comparison is small. Here, we first compare our method to the original SRe2(Yin et al., 2023) to observe the direct effect of our regularizer on the adversarial robustness of the trained model. We also compare with MTT (Cazenavette et al., 2022) and TESLA (Cui et al., 2023) on the same datasets to gain a further understanding on the differences in robustness between our method and other dataset distillation methods. We utilized the exact ConvNet architecture described in the papers of MTT and TESLA for their distillation and evaluation, as their performance on ResNet seems to be significantly lower.
@@
We evaluate all the methods on three distillation scales: 10 ipc, 50 ipc, and 100 ipc. We also employed a range of attacks to evaluate the robustness of the model, including PGD100 (Madry et al., 2017), Square (Andriushchenko et al., 2020), AutoAttack (Croce and Hein, 2020), CW (Carlini and Wagner, 2017), and MIM (Dong et al., 2017). This assortment includes both white-box and black-box attacks, providing a comprehensive evaluation of GUARD. For all adversarial attacks, with the exception of CW attack, we use the setting $\epsilon=1/255$. For CW specifically, we set the box constraint $c$ to $1e^{-5}$. Due to computational limits, we were not able to obtain results for MTT and TESLA with the 100 ipc setting on ImageNet, as well as the 100 ipc setting on ImageNet for all methods.
5.3 Results
@@
We provide a visualization of the distilled images generated by GUARD in Fig. 1, utilizing a distillation scale of 1 image per class among selected ImageNet classes. It can be seen that the images exhibit characteristics that resemble a blend of multiple objects within their assigned class, highlighting the method’s capacity to capture essential features.
The results of our experiment are detailed in Table 2. It can be observed that GUARD consistently outperforms both SRe2L and MTT in terms of robustness across various attacks.
Interestingly, we observe a notable increase in clean accuracy upon incorporating GUARD. While enhancing clean accuracy was not the primary goal of GUARD, this outcome aligns with its function as a regularizer, potentially aiding in model generalization. In the context of dataset distillation, where the goal is to distill essential features of the original dataset into a much smaller subset, improving the generalization is expected to have positive effects on the performance.
@@
GUARD
SRe2L
MTT
TESLA
ImageNette 10 None (Clean) 57.93
42.42
58.43
36.84
PGD100
23.87
4.76
39.85 28.10
Square
44.07
22.77
34.79
24.61
AutoAttack
19.69
4.99
33.72 24.48
CW
41.47
22.11
34.57
24.61
MIM
21.80
4.76
39.20 28.15
50 None (Clean) 80.86 80.15
59.69
36.21
PGD100
41.42
12.30
41.13
28.72
Square
72.81 61.50
36.72
25.34
AutoAttack
42.47
12.91
35.46
27.21
CW
58.67 53.42
36.54
29.01
MIM
43.23
12.43
41.69
30.12
100 None (Clean) 83.39 85.83
64.33
45.04
PGD100
57.50
31.65
44.89
33.98
Square
77.68
19.18
40.41
29.27
AutoAttack
64.84
17.93
39.46
28.99
CW
69.35 68.20
40.66
29.32
MIM
65.07
18.98
44.89
33.98
TinyImageNet 10 None (Clean) 37.00 33.18
8.14
14.06
PGD100
6.39
1.08
4.08
8.40
Square
19.53 15.85
2.48
6.31
AutoAttack
4.91
0.79
2.44
6.16
CW
8.40
3.24
2.50
6.26
MIM
6.51
1.10
4.08
8.40
50 None (Clean) 55.61 56.42
17.84
28.24
PGD100
15.63
0.27
5.62
12.12
Square
36.93 15.50
3.84
10.39
AutoAttack
13.84
0.16
3.52
10.01
CW
20.46 12.12
3.66
10.13
MIM
16.09
0.29
5.64
12.12
100 None (Clean) 60.13 59.30
29.16
30.48
PGD100
13.79
0.25
8.63
14.45
Square
37.06 17.74
7.29
12.02
AutoAttack
12.76
0.19
6.75
11.57
CW
20.05 14.02
6.93
11.57
MIM
14.35
0.24
8.63
14.45
ImageNet-1K 10 None (Clean) 27.25
21.30
PGD100
5.25 0.55
Square
17.88 18.02
AutoAttack
3.33 0.34
CW
7.68 3.21
MIM
5.23 0.51
50 None (Clean) 39.89 46.80
PGD100
9.77 0.59
Square
28.39 32.40
AutoAttack
7.03 0.47
CW
14.14 6.31
MIM
9.84 0.64
5.4 Ablation
In Section 4.2, Eq. 7 showed that the adversarial loss is upper-bounded by the normal loss, the gradient magnitude, and the curvature term. GUARD regularizes the curvature term while disregarding the gradient magnitude, which could theoretically reduce the upper bound of the loss as well. Here, we investigate the effect of regularizing gradient instead of curvature and present the results in Table 3. The results indicate that GUARD outperforms the gradient regularization alternatives, regardless of the regularization parameter.
Table 3: Accuracy on ImageNette of original SRe2L, GUARD, and gradient regularization on SRe2L with regularization parameters ($\lambda_{g}=10^{-4},10^{-3},10^{-2},0.1,1$). The best results among all methods are highlighted in bold.
@@ 5.4 Ablation
SRe2L GUARD $\lambda_{g}=10^{-4}$ $\lambda_{g}=10^{-3}$ $\lambda_{g}=10^{-2}$ $\lambda_{g}=0.1$ $\lambda_{g}=1$
1 None (Clean) 27.97 37.49 13.72 16.15 16.41 17.58 18.75
@@
6 Discussion
6.1 Robustness Guarantee
Due to the nature of dataset distillation, it is impossible to optimize the robustness of the final model with respect to the real dataset. Therefore, most approaches in this direction, including ours, have to optimize the adversarial loss of the model with respect to the distilled dataset. Unfortunately, there is always a distribution shift between the real and distilled datasets, which raises uncertainties about whether robustness on the distilled dataset will be effectively transferred when evaluated against the real dataset.
Nevertheless, our theoretical framework offers assurances regarding this concern. A comparison between Eq. 6 with Eq. 7 reveals that the bounds of adversarial loss for real data and distilled data differ only by $L\sigma$. For a well-distilled dataset, $\sigma$ should be relatively small. We have thus demonstrated that the disparity between minimizing adversarial loss on the distilled dataset and on the real dataset is confined to this constant. This conclusion of our theory allows future robust dataset distillation methods to exclusively enhance robustness with respect to the distilled dataset, without worrying if the robustness can transfer well to the real dataset.
6.2 Computational Overhead
The structure of robust dataset distillation, as outlined in Eq. 4, inherently presents a tri-level optimization challenge. Typically, addressing such a problem could entail employing complex tri-level optimization algorithms, resulting in significant computational demands. One example of this is the integration of adversarial training within the distillation framework, which necessitates an additional optimization loop for generating adversarial examples within each iteration.
However, GUARD’s approach, as detailed in Eq. 11, introduces an efficient alternative. GUARD’s regularization loss only requires an extra forward pass to compute the loss $\ell(\mathbf{x}+h\mathbf{z})$ within each iteration. Therefore, integrating GUARD’s regularizer into an existing method does not significantly increase the overall computational complexity, ensuring that the computational overhead remains minimal. This efficiency is particularly notable given the computationally intensive nature of tri-level optimization in robust dataset distillation. In Table 4, we present a comparison of the time per iteration required for a tri-level optimization algorithm, such as the one used for embedded adversarial training described in Section 3.2, against the time required for GUARD. The findings show that GUARD is much more computationally efficient and has a lower memory usage as well.
Table 4: Computation overhead of GUARD compared with embedded adversarial training. Experiments are performed on one NVIDIA A100 80GB PCIe GPU with batch size 32. We measure 5 times per iteration training time and report the average and standard deviation. Adv Training follows Section 3.2.

### v4

@@ Methods
Theoretical Bound of Robustness
Our aim is to create a method that allows us to efficiently and reliably introduce robustness into distilled datasets, thus we will start by exploring the theoretical connections between dataset distillation and adversarial robustness. Conveniently, previous works (Jetley, Lord, and Torr 2018; Fawzi et al. 2018) have studied the adversarial robustness of neural networks via the geometry of the loss landscape. Inspired by Moosavi-Dezfooli, Fawzi, and Frossard (2016), here we find connections between standard training procedures and dataset distillation to provide a theoretical bound for the adversarial loss of models trained with distilled datasets.
@@
$\begin{gathered}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\tilde{\ell}_{\rho}^{adv}(\mathbf{x})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|{\nabla\ell(\mathbf{x})}\|+\\ \frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x}),\end{gathered}$ (8)
where $\lambda_{1}$ is the largest eigenvalue of the Hessian matrix $\mathbf{H(\ell(x))}$. Then, we have the proposition:
@@
~$\begin{gathered}\tilde{\ell}_{\rho}^{adv}(\mathbf{x}^{\prime})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|\nabla\ell(\mathbf{x})\|+\\ \frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x})+L\sigma.\end{gathered}$ (11)
Given the assumption of convexity in the loss function, we can further observe that in a convex landscape the gradient magnitude tends to be lower near the optimal points. Therefore, in the context of a convex loss function and a well-distilled dataset, the gradient term $\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|\nabla\ell(\mathbf{x})\|$ contribute insignificantly to RHS of Eq. 11. This insignificance is amplified by the presence of the curvature term, $\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x})$, which provides a sufficient descriptor of the loss landscape under our assumptions. Hence, it is reasonable to simplify the expression by omitting the gradient term, resulting in a focus on the curvature term, which is more representative of the convexity assumption and the characteristics of a well-distilled dataset. The revised expression would then be:
@@
In the appendix, we give a more thorough proof of the proposition and discuss the validity of some of the assumptions made. In the Experiments section, we also show results from an ablation study to demonstrate the empirical effects of some of these assumptions.
Geometric Regularization for Adversarially Robust Dataset
Based on our theoretical discussion, we propose a method, GUARD (Geometric Regularization for Adversarially Robust Dataset). Since the theorem suggests that the upper bound of the adversarial loss is mainly determined by the curvature of the loss function, we modify the distillation process so that the trained model has a loss function with a low curvature with respect to real data.
Reducing $\lambda_{1}$ in Eq. 12 requires computing the Hessian matrix to get the largest eigenvalue $\lambda_{1}$, which is quite computationally expensive. Here we find an efficient approximation of it. Let $\mathbf{v_{1}}$ be the unit eigenvector corresponding to $\lambda_{1}$, then the Hessian-vector product is
@@
Previous works (Fawzi et al. 2018; Jetley, Lord, and Torr 2018; Moosavi-Dezfooli et al. 2019) have empirically shown that the direction of the gradient has a large cosine similarity with the direction of $\mathbf{v_{1}}$ in the input space of neural networks. Instead of calculating $\mathbf{v_{1}}$ directly, it is more efficient to take the gradient direction as a surrogate of $\mathbf{v_{1}}$ to perturb the input $\mathbf{x}$. So we replace the $\mathbf{v}_{1}$ above with the normalized gradient $\mathbf{z}=\frac{\nabla\ell(\mathbf{x}))}{\|\nabla\ell(\mathbf{x}))\|}$, and define the regularized loss $\ell_{R}$ to encourage linearity in the input space:
@@
Experiments
Experiment Settings
@@
For a systematic evaluation of our method, we investigate the top-$1$ classification accuracy of models trained on data distilled from three commonly-used datasets in this area: ImageNette (Howard 2018), Tiny ImageNet (Le and Yang 2015), and ImageNet-1K (Deng et al. 2009). ImageNette is a subset of ImageNet-1K containing 10 easy-to-classify classes. Tiny ImageNet is a scaled-down subset of ImageNet-1K, containing 200 classes and 100,000 downsized 64x64 images. We trained networks using the distilled datasets and subsequently evaluated the network’s performance on the validation split of the original datasets (because none of these datasets have a test split with publicly available labels). For consistency in our experiments across all datasets, we used the standard ResNet18 architecture (He et al. 2016) to synthesize the distilled datasets and evaluate their performance.
During the squeeze step of the distillation process, we trained the model on the original dataset over 50 epochs using a learning rate of 0.025. Based on preliminary experiments, we determined that the settings $h=3$ and $\lambda=100$ provide an optimal configuration for our regularizer. In the recover step, we performed 2000 iterations to synthesize the images and run 300 epochs to generate the soft labels to obtain the full distilled dataset. In the evaluation phase, we trained a ResNet18 model on the distilled dataset for 300 epochs, before assessing it on the validation split of the original dataset.
Comparison with Other Methods
@@
As of now, there is only a small number of dataset distillation methods that can achieve good performance on ImageNet-level datasets, therefore our choices for comparison is small. Here, we first compare our method to the original SRe2(Yin, Xing, and Shen 2023) to observe the direct effect of our regularizer on the adversarial robustness of the trained model. We also compare with MTT (Cazenavette et al. 2022) and TESLA (Cui et al. 2023) on the same datasets to gain a further understanding on the differences in robustness between our method and other dataset distillation methods. We utilized the exact ConvNet architecture described in the papers of MTT and TESLA for their distillation and evaluation, as their performance on ResNet seems to be significantly lower.
@@
We evaluate all the methods on three distillation scales: 10 IPC, 50 IPC, and 100 IPC. We also employed a range of attacks to evaluate the robustness of the model, including PGD100 (Madry et al. 2017), Square (Andriushchenko et al. 2020), AutoAttack (Croce and Hein 2020), CW (Carlini and Wagner 2017), and MIM (Dong et al. 2017). This assortment includes both white-box and black-box attacks, providing a comprehensive evaluation of GUARD. For all adversarial attacks, with the exception of CW attack, we use the setting $\epsilon=1/255$. For CW specifically, we set the box constraint $c$ to $10^{-5}$. Due to computational limits, we were not able to obtain results for MTT and TESLA with the 100 IPC setting on ImageNet, as well as the 100 IPC setting on ImageNet for all methods.
Results
@@
The results are detailed in Table 2. It can be observed that GUARD consistently outperforms both SRe2L and MTT in terms of robustness across various attacks. Interestingly, we observe an increase in clean accuracy upon incorporating GUARD across various settings. While enhancing clean accuracy was not the primary goal of GUARD, this outcome aligns with its function as a regularizer, potentially aiding in model generalization. In the context of dataset distillation, where the goal is to distill essential features of the original dataset into a smaller subset, improving the generalization is expected to have positive effects on the performance. We also provide a visualization of the distilled images generated by GUARD in Figure 1, utilizing a distillation scale of 1 image per class among selected ImageNet classes. It can be seen that the images exhibit characteristics that resemble a blend of multiple objects within their assigned class, highlighting the method’s capacity to capture essential features.
@@
GUARD SRe2L MTT TESLA
ImageNette 10 None (Clean) 57.93 42.42 58.43 36.84
PGD100 23.87 4.76 39.85 28.10
Square 44.07 22.77 34.79 24.61
AutoAttack 19.69 4.99 33.72 24.48
CW 41.47 22.11 34.57 24.61
MIM 21.80 4.76 39.20 28.15
50 None (Clean) 80.86 80.15 59.69 36.21
PGD100 41.42 12.30 41.13 28.72
Square 72.81 61.50 36.72 25.34
AutoAttack 42.47 12.91 35.46 27.21
CW 58.67 53.42 36.54 29.01
MIM 43.23 12.43 41.69 30.12
100 None (Clean) 83.39 85.83 64.33 45.04
PGD100 57.50 31.65 44.89 33.98
Square 77.68 19.18 40.41 29.27
AutoAttack 64.84 17.93 39.46 28.99
CW 69.35 68.20 40.66 29.32
MIM 65.07 18.98 44.89 33.98
TinyImageNet 10 None (Clean) 37.00 33.18 8.14 14.06
PGD100 6.39 1.08 4.08 8.40
Square 19.53 15.85 2.48 6.31
AutoAttack 4.91 0.79 2.44 6.16
CW 8.40 3.24 2.50 6.26
MIM 6.51 1.10 4.08 8.40
50 None (Clean) 55.61 56.42 17.84 28.24
PGD100 15.63 0.27 5.62 12.12
Square 36.93 15.50 3.84 10.39
AutoAttack 13.84 0.16 3.52 10.01
CW 20.46 12.12 3.66 10.13
MIM 16.09 0.29 5.64 12.12
100 None (Clean) 60.13 59.30 29.16 30.48
PGD100 13.79 0.25 8.63 14.45
Square 37.06 17.74 7.29 12.02
AutoAttack 12.76 0.19 6.75 11.57
CW 20.05 14.02 6.93 11.57
MIM 14.35 0.24 8.63 14.45
ImageNet-1K 10 None (Clean) 27.25 21.30 - -
PGD100 5.25 0.55 - -
Square 17.88 18.02 - -
AutoAttack 3.33 0.34 - -
CW 7.68 3.21 - -
MIM 5.23 0.51 - -
50 None (Clean) 39.89 46.80 - -
PGD100 9.77 0.59 - -
Square 28.39 32.40 - -
AutoAttack 7.03 0.47 - -
CW 14.14 6.31 - -
MIM 9.84 0.64 - -
Ablation Study on Gradient Regularization
Eq. 11 showed that the adversarial loss is upper-bounded by the normal loss, the gradient magnitude, and the curvature term. GUARD regularizes the curvature term while disregarding the gradient magnitude, which could theoretically reduce the upper bound of the loss as well. Here, we investigate the effect of regularizing gradient instead of curvature and present the results in Table 3. The results indicate that GUARD outperforms the gradient regularization alternatives, regardless of the regularization parameter.
Table 3: Accuracy on ImageNette of original SRe2L, GUARD, and gradient regularization on SRe2L with regularization parameters ($\lambda_{g}=10^{-4},10^{-3},10^{-2},0.1,1$, where $\lambda_{g}$ is omitted in table columns for brevity). AA stands for AutoAttack. The best results among all methods are highlighted in bold.
@@
SRe2L GUARD $10^{-4}$ $10^{-3}$ $10^{-2}$ $0.1$ $1$
1 (Clean) 27.97 37.49 13.72 16.15 16.41 17.58 18.75
@@
Discussion
Robustness Guarantee
Due to the nature of dataset distillation, it is impossible to optimize the robustness of the final model with respect to the real dataset. Therefore, most approaches in this direction, including ours, have to optimize the adversarial loss of the model with respect to the distilled dataset. Unfortunately, there is always a distribution shift between the real and distilled datasets, which raises uncertainties about whether robustness on the distilled dataset will be effectively transferred when evaluated against the real dataset. Nevertheless, our theoretical framework offers assurances regarding this concern. A comparison between Eq. 8 and Eq. 11 reveals that the bounds of adversarial loss for real data and distilled data differ only by $L\sigma$. For a well-distilled dataset, $\sigma$ should be relatively small. We have thus demonstrated that the disparity between minimizing adversarial loss on the distilled dataset and on the real dataset is confined to this constant. This conclusion of our theory allows future robust dataset distillation methods to exclusively enhance robustness with respect to the distilled dataset, without worrying if the robustness can transfer well to the real dataset.
Computational Overhead
The structure of robust dataset distillation, as outlined in Eq. 4, inherently presents a tri-level optimization challenge. Typically, addressing such a problem could entail employing complex tri-level optimization algorithms, resulting in significant computational demands. One example of this is the integration of adversarial training within the distillation framework, which necessitates an additional optimization loop for generating adversarial examples within each iteration. However, GUARD’s approach, as detailed in Eq. 15, introduces an efficient alternative. GUARD’s regularization loss only requires an extra forward pass to compute the loss $\ell(\mathbf{x}+h\mathbf{z})$ within each iteration. Therefore, integrating GUARD’s regularizer into an existing method does not significantly increase the overall computational complexity, ensuring that the computational overhead remains minimal. This efficiency is particularly notable given the computationally intensive nature of tri-level optimization in robust dataset distillation. In Table 4, we present a comparison of the time per iteration required for a tri-level optimization algorithm, such as the one used for embedded adversarial training, against the time required for GUARD. The findings show that GUARD is much more computationally efficient and has a lower memory usage as well.
Table 4: Computation overhead of GUARD compared with embedded adversarial training. Experiments are performed on one NVIDIA A100 80GB PCIe GPU with batch size 32. We measure 5 times per iteration training time and report the average and standard deviation.